OpenStudy (anonymous):

expand f(x)=[x+sqr(x)]/[(1-x)^3] as a power series and use it to find the sum of sigma n^2/2^n

OpenStudy (amistre64):

I think I would try to get the bottom rationalized

OpenStudy (amistre64):

nah, expanded; since ^3 aint a radical :)

OpenStudy (amistre64):

then divide it out

OpenStudy (amistre64):

1 - x - x^3 i beleive is the bottom expanded

OpenStudy (amistre64):

er, +x

OpenStudy (amistre64):

x - x^2 -------------------- 1 + x - x^3 ) x + sqrt(x) (x+x^2-x^4) ------------- -x^2+x^4+sqrt(x) (-x^2-x^3+x^5) ----------------- x^3+x^4 +x^5 +sqrt(x) hmmm, should we let x^1/2 come before x ? or continue as planned?

OpenStudy (anonymous):

hi this is a power series question

OpenStudy (amistre64):

if anyting we could split the numberator

OpenStudy (amistre64):

yes, and this is one way to build a power series .... or do you have something more specific in mind?

OpenStudy (anonymous):

yah...derivative of 1/1-x gives 1/(1-x)2 and again differentiating gives 1/(1-x)^3 like this way we can solve

OpenStudy (amistre64):

then determine the derivatives of f(x) till you see a pattern emerget hat can notated in summation

OpenStudy (anonymous):

can u answer my second part of the question

OpenStudy (amistre64):

let me consider it .... see if i can make sense of what is being asked :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

when I split up the numerator into x and sqrt(x) I get similar expansions: \[\frac{x}{(1-x)^3}=x+3x^2+6x^3+10x^3+15x^4+21x^5...\] \[\frac{x^{1/2}}{(1-x)^3}=x^{1/2}+3x^{3/2}+6x^{5/2}+10x^{7/2}+15x^{9/2}+21x^{11/2}...\]

OpenStudy (amistre64):

but I aint sure how to use it to find the second part